# Robust Understanding of Human-Robot Social Interactions through Multimodal Distillation

### v1

@@ 3.5. Input Corruption
In real-world applications, the social understanding model must not only work under the computational constraints of mobile platforms but also remain robust to incomplete and noisy inputs. To simulate such conditions, we deliberately introduce corruption in both spatial and temporal dimensions during training. In the spatial domain, a fixed number of body pose keypoints are randomly selected and corrupted by either setting their coordinate values and confidence scores to zero or replacing the coordinates with random values within the normalized range. This mimics localized body occlusions or pose estimation errors. In the temporal domain, we randomly drop a fixed number of frames by setting all keypoint values in those frames to zero or normalized random noise, simulating transient global occlusions. These corruptions force the student model to learn robust representations and maintain performance under incomplete or noisy input during knowledge distillation.
@@ 4.1. Data
Figure 3. We standardized the input frame rate of two datasets to 10 FPS and limited the observation window to the first second of social interaction.
To evaluate the performance and robustness of our models in forecasting users’ social intents and behaviours from an egocentric perspective, we use two datasets of human-robot social interactions recorded from the robot’s perspective. Examples of datasets are shown in Fig. 3.
JPL-Social (Bian et al., 2025) is an extension of the JPL-Interaction datasets (Ryoo and Matthies, 2013; Ryoo et al., 2015), designed for human-robot interaction research, with a focus on users’ social actions within the central visual field. The original dataset captures 8 social actions performed by the main user: handshake, hug, pet, wave, punch, throw, point, and leave. JPL-Social extends this by annotating additional users in the field of view during multi-person scenarios, introducing two additional social actions: gaze and no response. Furthermore, JPL-Social provides labels for social intent (categorized as interested, interacting, or not interested) and attitude (positive or negative) for each user. The dataset comprises 290 interaction sequences involving 16 participants in diverse environmental settings.
HARPER (Avogaro et al., 2024) is a dataset capturing 10 human-robot social actions performed by 17 users, recorded from multiple perspectives using five cameras mounted on a quadrupedal robot. For our research, we use 6 social actions specifically relevant to robot-oriented interactions: crash, walk-avoid, touch, walk-stop, punch, and kick, so there are 260 interaction sequences. In addition, HARPER includes variations in viewing angle and resolution, increasing the difficulty of understanding users’ social signals. To ensure consistency in social representation, we manually annotated HARPER with the same social intent and attitude labels as JPL-Social.
@@ 4.2. Experiment Setting
We set the Frames Per Second (FPS) of the videos to 10 and limited the observation window to the first second of social interaction, which starts when the user approaches the social agent, so the inputs have 10 frames. To expand the training set, we randomly cropped video frames and re-extracted pose features, facial expressions, hand gestures, and gaze information. After horizontal flipping, we augmented number of videos 20 times for richer training data. We train and test the models on the two datasets separately. Given our focus on model efficiency, we report both the number of model parameters and the number of floating point operations (FLOPs) for each input.
@@ 4.3. Performance of the Teacher Model
Model Params FLOPs JPL-Social HARPER
(M) (M) Intent Acc. Attitude Acc. Action Acc. Intent Acc. Attitude Acc. Action Acc.
ST-GCN+ (Yan et al., 2018) 43.86 1924.87 86.90 76.19 71.43 86.54 73.08 78.85
ST-TR+ (Plizzari et al., 2021) 58.48 4724.63 79.76 59.52 48.81 75.00 65.38 59.61
MS-G3D+ (Liu et al., 2020) 48.82 5241.97 88.10 80.95 76.19 90.38 78.85 80.77
SocialEgoNet+ (Bian et al., 2025) 37.78 422.35 86.90 77.38 71.43 86.54 75.00 78.85
SocialC3D (ours) 48.49 7902.19 92.85 88.10 82.14 96.15 82.69 88.46
SocialEgoMobile (ours) 0.43 4.23 82.14 71.43 67.86 69.23 44.23 51.92
Table 1. Comparison of SocialC3D and SocialEgoMobile with state-of-the-art methods on the JPL-Social and HARPER. SocialEgoMobile relies solely on clean body pose features as input. ’+’ indicates that the model uses raw image and gaze information.
As the teacher model, we compare SocialC3D with several state-of-the-art methods. Although originally designed for body pose skeleton graphs, these methods can also process facial and hand keypoint graphs. To ensure a fair comparison, we incorporated the raw image and gaze information channels from SocialC3D into all teacher models, aligning the available modalities across teacher models. Each modality channel in the teacher models is trained independently first and then during joint fine-tuning, outputs of all the channels are fused before being passed to the chain hierarchical classifier. To enhance body language understanding, the body pose and raw image channels of all teacher models are initialized with pre-trained weights on Kinetics 400 (Kay et al., 2017).
@@
As shown in Tab le1, we compare SocialC3D with SOTA methods on the JPL-Social and HARPER datasets. SocialC3D consistently achieves superior performance on three downstream social tasks in both datasets. Compared with the second-best model, MS-G3D (Liu et al., 2020), SocialC3D achieves an average accuracy improvement of 5.85% on JPL-Social and 5.77% on HARPER, which highlights the effectiveness of SocialC3D in understanding social cues and forecasting users’ social intentions and behaviours from an egocentric perspective. We attribute this improvement to two key factors. First, instead of directly using pose features, converting them into heatmaps enhances robustness to pose estimation errors, which is particularly pronounced in egocentric settings due to narrow and jittery fields of view. Second, our experiments focus on the initial 10 frames of social interactions. 3D CNNs are suitable for capturing local temporal dynamics, making them more appropriate for our experiment setting. In contrast, models like ST-GCN (Yan et al., 2018) and MS-G3D (Liu et al., 2020) rely on complete and accurate pose features, making them more vulnerable to missing joints and estimation noise. ST-TR (Plizzari et al., 2021), while effective in modelling long-term dependencies due to using transformers (Vaswani et al., 2017), struggles with short sequences and lacks mechanisms to handle input degradation. Although SocialEgoNet (Bian et al., 2025) is designed to capture users’ non-verbal social cues from an egocentric perspective, its lightweight architecture limits its capacity to extract fine-grained information independently.
@@
In addition, we report the performance of SocialEgoMobile using only body pose to compare its accuracy and computational efficiency against the teacher models. Compared to SocialC3D, SocialEgoMobile’s average accuracy of downstream tasks on JPL-Social was 13. 89% lower and 33.97% lower on HARPER. The larger performance gap on HARPER is attributed to its recording setup, where the robot dog has a lower vantage point and fewer frames with full-body visibility, resulting in noisier and more incomplete pose inputs, as shown in Fig.3. In terms of computational efficiency, compared to SocialC3D, SocialEgoMobile only needs 0.98% of the model parameters and 0.54‰of FLOPs. These results show SocialEgoMobile is computationally lightweight enough for social understanding on mobile platforms with limited resources.
@@ 4.4. Robustness Analysis on the Student Model
Figure 4. Knowledge distillation (KD) consistently improves the performance of the student model, SocialEgoMobile, under Individual and simultaneous spatio-temporal corruption on all three downstream tasks, interaction intent, attitude, and social action forecast. Improvements on downstream task accuracy through distillation are labelled.
To address the practical challenges faced by mobile socially intelligent agents from an egocentric perspective, such as occlusion of user body parts and pose estimation error, we deliberately introduced random spatio-temporal corruption into the input of the student model SocialEgoMobile, as mention in Section 3.5. The goal is to improve robustness by using the rich and clean knowledge extracted from the intermediate social representations of the teacher model SocialC3D, to supervise the learning of the student model with corrupted inputs. During the robustness tests, student models were trained for 70 epochs. The first 30 epochs are warm-up phase, where the corruption rate was gradually increased to the final value to prevent unstable training due to excessive information loss.
Fig.4 compares SocialEgoMobile’s performance under two settings: independent learning and knowledge distillation. Under independent learning, performance degrades as corruption increases. When 30% of joints and frames is corrupted, $30\%\oplus 30\%=51\%$ of the information in the input was corrupted. The average accuracy across three downstream tasks drops by 8.85%, demonstrating the increased difficulty and robustness demands. In contrast, with knowledge distillation, SocialEgoMobile consistently outperforms the independently trained counterpart across all input settings. When 30% of joints and frames is corrupted, SocialEgoMobile trained via distillation surpasses the performance of the independently trained model using clean input. This suggests that knowledge distillation effectively transfers social knowledge from the teacher model, which incorporates additional modalities and complete, clean inputs, to the student model. So it can more accurately capture valuable social cues even when facing corrupted inputs. The performance gains are particularly significant for the user attitude and social action prediction tasks, which rely more on detailed and accurate input.
Whether trained independently or via knowledge distillation, SocialEgoMobile is more sensitive to the corruption of spatial information than temporal information. This is because temporal patterns in human behaviour tend to be more regular and predictable, allowing models to more effectively compensate for missing frames and filter out noise. In contrast, spatial patterns of pose graphs are more complex, such as body structure and human kinematics.
@@ 4.5. Ablation Experiments
Figure 5. Performance comparison on JPL-Social and HARPER of SocialC3D using different input modalities, and the impact of corresponding SocialC3D-based Knowledge Distillation on SocialEgoMobile (SEM) when 30% of spatio-temporal information are corrupted. Accuracy drops due to missing modalities are labelled.
Contribution of Mutli-Modalities.  We incorporate body pose, raw images, face, hand gestures, and gaze as input modalities for the teacher model SocialC3D, enabling it to comprehensively capture social cues from users’ non-verbal behaviours across multiple perspectives. Body pose and raw images are the foundational modalities and prior work has demonstrated their effectiveness and complementarity in understanding body movements (Duan et al., 2022). In contrast, facial expressions, hand gestures, and gaze focus on localized regions of the body, providing fine-grained cues essential for capturing nuanced social signals.
Fig.5 illustrates SocialC3D’s performance across three downstream tasks with full modality and with each of the three localized modalities ablated. The Attitude and Intent tasks show performance degradation when facial information is removed. Facial orientation offers cues about a user’s attention, while expressions convey emotional states that differentiate social behaviours with similar body movements in early stages, such as ‘throw’ and ‘wave hands’. The Action task is more sensitive to the absence of hand gestures, as detailed hand pose information helps the model to distinguish between fine-grained actions, such as the differences in finger closure for ‘touch’ and ‘handshake’. For the Intent task, the lack of gaze information leads to the greatest drop in accuracy, since gaze direction directly reflects user focus and intention. While facial orientation may offer partial substitutes, it cannot fully replicate the precision of direct gaze signals. Through knowledge distillation, the impact of missing modalities observed in the teacher model is transferred to the student model, as shown in Fig.5. These results confirm that the proposed knowledge distillation framework successfully transfers the integrated multimodal social knowledge from the teacher model to the unimodal student model.
@@
Distillation Methods.  In Table 2, we compare the performance improvements brought by different knowledge distillation methods to the student model, SocialEgoMobile, when 30% of joints and frames are corrupted. Distillation based on intermediate features outperforms soft label distillation, as the student model in our experiments must learn to integrate information of additional modalities from high-level social representation provided by the teacher model and compensate for the impact of missing or noisy inputs. Learning only from the teacher’s output distributions is insufficient. Among the intermediate feature-based methods, InfoNCE (Oord et al., 2018) achieves the best performance. By maximizing mutual information between the teacher and student model’s intermediate social representations, InfoNCE encourages the student to learn semantically aligned features, rather than enforcing exact numerical matching at lower levels. This is suitable for our experiment setting, where the teacher and student operate on different input modalities and the student receives incomplete and noisy inputs, making low-level alignment impractical. InfoNCE promotes the alignment of high-level semantic information, leading to better noise tolerance and generalization.
Model Architecture JPL-Social
Intent F1 Attitude F1 Action F1
SocialEgoMobile 88.10 80.95 75.00
w/ 3 layers GAT 84.52 76.19 69.05
w/ 1 layers GAT 83.33 69.05 67.86
w/ 2 layer GCN (Kipf and Welling, 2017) 85.71 71.43 67.86
w/ 2 layers Bi-LSTM 88.10 82.14 75.00
w/ 1 layer TCN (Lea et al., 2017) 83.33 69.05 65.48
w/ 1 layer Transformer (Vaswani et al., 2017) 77.38 66.67 54.76
@@
Table 3. Performance Comparison of SocialEgoMobile with Different Backbone Models on JPL-Social when 30% joints and frames are corrupted. The best results are shown in bold and the second best results are underlined.
@@
Student Model Architecture.  To enhance the robustness and efficiency of processing noisy pose feature sequences, we designed the student model SocialEgoMobile with a two-layer GAT (Veličković et al., 2018) followed by a one-layer Bi-LSTM (Graves and Schmidhuber, 2005). Table 3 presents a comparison of different model configurations under the knowledge distillation with corrupted inputs. In the spatial dimension, a two-layer GAT achieves a good trade-off between expressiveness and stability. Fewer layers limit the receptive field, reducing the model’s ability to capture global graph structural patterns, while deeper GATs tend to propagate noise more widely. Moreover, GAT’s adaptive attention mechanism enables selective aggregation of useful information from neighbouring nodes, offering greater resilience to noisy or missing joints compared to standard GCNs (Kipf and Welling, 2017). In the temporal dimension, a single-layer Bi-LSTM provides sufficient modelling capacity for short pose feature sequences, while deeper Bi-LSTMs introduce additional parameters with marginal performance improvement. Compared to TCNs (Lea et al., 2017) and Transformers (Vaswani et al., 2017), Bi-LSTM is better suited for sequential data with inherent temporal dependencies and benefits from its gating mechanism, which improves robustness to sparsity and temporal discontinuity. Overall, the combination of two GAT layers and one Bi-LSTM in SocialEgoMobile offers an effective balance between expressiveness, robustness, and computational efficiency.

### v2

@@ 3.5. Input Corruption
In real-world applications, social understanding models must work under the computational constraints of mobile platforms and be robust to incomplete and noisy inputs. To simulate such conditions, we deliberately introduce corruption in spatial and temporal dimensions. In the spatial domain, a fixed number of body pose keypoints are randomly corrupted by setting their coordinate values and confidence scores to zero or replacing the coordinates with normalized random values. This mimics localized body occlusions and pose estimation errors. In the temporal domain, we randomly corrupt frames by setting all keypoint values to zero or applying normalised random noise, simulating transient full-body occlusions. These corruptions force the student model to learn robust representations and maintain performance under incomplete and noisy input.
@@ 4.1. Data
Figure 3. We set the FPS of two datasets to 10 and limited the observation window to the first second of social interactions.
To evaluate the performance and robustness of our models in egocentric social understanding, we use two datasets of human-robot social interactions recorded from the robot’s perspective. Examples of the datasets are shown in Fig. 3.
JPL-Social (Bian et al., 2025) is an extension of the JPL-Interaction datasets (Ryoo and Matthies, 2013; Ryoo et al., 2015), focused on the user’s social actions. The original dataset captures 8 social actions performed by the users in the centre of the screen: handshake, hug, pet, wave, punch, throw, point, and leave. JPL-Social extends this by annotating additional users in the background during multi-person scenarios, introducing two additional action label: gaze and no response. Furthermore, JPL-Social provides annotations for social intent (categorized as interacting, interested and not interested) and attitude (positive or negative) for each user. The dataset comprises 290 interaction sequences involving 16 participants in diverse environmental settings.
HARPER (Avogaro et al., 2024) is a dataset recorded from multiple perspectives using five cameras mounted on a quadrupedal robot. We use 6 social actions relevant to robot-oriented interactions: crash, walk-avoid, touch, walk-stop, punch and kick, so there are 260 sequences. In addition, HARPER includes variations in viewing angle and resolution, increasing the difficulty of understanding users’ behaviours. Following the protocol of JPL-Social (Bian et al., 2025), we manually annotate HARPER for the intent and attitude prediction tasks.
@@ 4.2. Experiment Setting
We set the Frames Per Second (FPS) of the videos to 10 and limited the observation window to the first second of social interaction, which starts when the user approaches the social agent, so the inputs have 10 frames. To expand the training set, we randomly cropped and horizontally flipped video frames, augmenting the number of videos 20 times for richer training data. We train and test the models on the two datasets separately. Given our focus on model efficiency, we report the number of model parameters and the average latency for each input. All experiments were conducted on an NVIDIA RTX 4070 GPU.
All models are trained for 40 epochs with a batch size of 96. Training is early stopped if the average performance across the three subtasks does not improve for 5 epochs. Each modality channel in the SocialC3D is trained independently first and then jointly fine-tuning. To enhance performance, the body pose and raw image channels of all teacher models are initialized with pre-trained weights on Kinetics 400 (Kay et al., 2017). SocialC3D is optimized using SGD (Sutskever et al., 2013), while SocialEgoMobile uses the Adam optimizer (Kingma and Ba, 2015). During knowledge distillation, the total loss is computed as a weighted sum of the classification loss and the distillation loss, with weights of 0.8 and 0.2. The temperature for the InfoNCE (Oord et al., 2018) is set to 0.1.
@@ 4.3. Performance of the Teacher Model
Table 1. Comparison of SocialC3D and SocialEgoMobile with state-of-the-art. ’+’ indicates that the model uses additional modalities same as SocialC3D. $\Delta_{1}$ denotes the time to extract whole-body pose and gaze features from a single frame, which is 4.96 ms under our experimental setup. $\Delta_{2}$ refers to the extraction time for body pose features, which is 3.06 ms.
Model Params Latency JPL-Social HARPER
Intent Attitude Action Intent Attitude Action
(M) $\downarrow$ (ms) $\downarrow$ Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
ST-GCN+ (Yan et al., 2018) 43.86 $\Delta_{1}$ + 10.31 86.90 85.93 76.19 74.56 71.43 65.23 86.54 84.02 73.08 74.68 78.85 77.18
ST-TR+ (Plizzari et al., 2021) 58.48 $\Delta_{1}$ + 14.33 79.76 77.82 59.52 58.87 48.81 46.76 75.00 66.37 65.38 58.91 59.61 39.63
MS-G3D+ (Liu et al., 2020) 48.82 $\Delta_{1}$ + 13.74 88.10 88.11 80.95 78.12 76.19 74.59 90.38 88.11 78.85 76.52 80.77 74.00
SocialEgoNet+ (Bian et al., 2025) 37.78 $\Delta_{1}$ + 10.05 86.90 86.54 77.38 75.68 71.43 68.93 86.54 77.29 75.00 73.72 78.85 71.78
SocialC3D 48.49 $\Delta_{1}$ + 22.34 92.85 91.25 88.10 88.91 82.14 78.09 96.15 96.14 82.69 88.28 88.46 88.35
SocialEgoMobile 0.43 $\Delta_{2}$ + 0.19 86.90 85.26 71.43 68.61 67.86 62.63 69.23 66.92 38.46 42.07 51.92 46.56
@@
As the teacher model, we compare SocialC3D with several state-of-the-art methods. Although originally designed for pose skeletons, these methods can also process facial and hand keypoints. To ensure a fair comparison, we incorporated the raw image and gaze information channels from SocialC3D into all teacher models, aligning the modalities across teacher models. In Table 1, SocialC3D consistently achieves superior performance on three downstream social tasks on both datasets. Compared with the second-best model, MS-G3D (Liu et al., 2020), SocialC3D achieves an average accuracy improvement of 5.85% on JPL-Social and 5.77% on HARPER, which highlights the effectiveness of SocialC3D in understanding social cues and forecasting users’ social intentions from an egocentric perspective. We attribute this improvement to two key factors. First, instead of directly using pose features, converting them into heatmaps enhances robustness to pose estimation errors, which is particularly pronounced in egocentric settings due to narrow and jittery fields of view. Second, our experiments focus on the first 10 frames of social interactions. 3D CNNs are suitable for capturing short-term temporal dynamics, making them more appropriate for our experiment setting. In contrast, models like ST-GCN (Yan et al., 2018) and MS-G3D (Liu et al., 2020) rely on complete and accurate pose features, making them more vulnerable to missing joints and estimation errors. ST-TR (Plizzari et al., 2021), while effective in modelling long-term dependencies due to using transformers (Vaswani et al., 2017), struggles with short sequences. Although SocialEgoNet (Bian et al., 2025) is designed to capture users’ non-verbal social cues from an egocentric perspective, its lightweight architecture limits its capacity to extract fine-grained information.
@@
In addition, we report the performance of SocialEgoMobile using only body pose to compare its accuracy and computational efficiency against the teacher models. Compared to SocialC3D, SocialEgoMobile’s average accuracy of downstream tasks was 13.89% lower on JPL-Social and 33.97% lower on HARPER. The larger performance gap on HARPER is attributed to its recording setup, where the robot dog has a lower viewing angle and fewer frames with full-body visibility, resulting in noisier and more incomplete pose inputs, as shown in Fig.3. In such conditions, high-level features such as body pose become less reliable. These features rely on per-frame feature estimation methods, which lack temporal context and will produce unstable outputs, leading to fragmented feature sequences and degraded model understanding. In contrast, raw image sequences provide a temporally coherent context. Thus, teacher models with image inputs can maintain high performance on HARPER, while SocialEgoMobile, relying solely on body pose, suffers from significant degradation. In terms of computational efficiency, we report the number of parameters and the inference latency for all methods. Since the pose and gaze estimation algorithms can run at 10 FPS, the reported latency only includes the model inference time and the feature extraction time for the last frame. SocialEgoMobile only needs 0.43 M model parameters (0.98% of SocialC3D) and 3.25 ms latency (11.9% of SocialC3D). These results show that SocialEgoMobile is lightweight enough for real-time social understanding on mobile platforms with limited resources.
@@ 4.4. Robustness Analysis on the Student Model
Figure 4. Knowledge distillation (KD) consistently improves the performance of the student model, SocialEgoMobile, under spatio-temporal corruption on all three downstream tasks, interaction intent, attitude, and social action forecast. Improvements on downstream task accuracy through distillation are labelled.
To address the practical challenges faced by mobile socially intelligent agents from an egocentric perspective, such as occlusion of user body parts and pose estimation error, we deliberately introduced random spatiotemporal corruption into the input of the student model SocialEgoMobile, as mentioned in Section 3.5. The goal is to improve robustness by using the multimodal and uncorrupted knowledge extracted from the teacher model, SocialC3D, to supervise the learning of the student model with corrupted inputs. During the robustness experiments, student models were trained for 70 epochs. The first 30 epochs are the warm-up phase, where the corruption rate is gradually increased to the target value to prevent unstable training due to excessive information loss.
Figure 5. Comparison of SocialC3D using different input modalities and the its impact on SocialEgoMobile (SEM). Accuracy drops due to missing modalities are labelled.
Fig.4 compares SocialEgoMobile’s performance under independent learning and knowledge distillation. Under independent learning, performance degrades as corruption increases. When 30% of joints and frames are corrupted, $30\%\oplus 30\%=51\%$ of the information in the input was corrupted. This ratio maintains a balance of noise and useful information. The average accuracy across three subtasks drops by 8.85%, demonstrating the increased difficulty and robustness demands. In contrast, with knowledge distillation, SocialEgoMobile consistently outperforms the independently trained counterpart across all input settings. When 30% of joints and 30% of frames are corrupted, SocialEgoMobile trained via distillation surpasses the performance of the independently trained one without input corruption. This suggests that knowledge distillation effectively transfers social knowledge from the teacher model, which incorporates additional modalities and complete inputs to the student model. So it can more accurately capture valuable social cues even when facing corrupted inputs. The performance gains are particularly significant for the user attitude and social action prediction tasks, which rely more on detailed and accurate input.
Whether trained independently or via knowledge distillation, SocialEgoMobile is more sensitive to the corruption of spatial information than temporal information. This is because temporal patterns in human behaviour tend to be more regular and predictable, allowing models to more effectively compensate for missing frames and filter out noise. In contrast, spatial patterns of pose skeletons are more complex, such as body structure and human kinematics.
@@ 4.5. Ablation Experiments
Contribution of Mutli-Modalities.  We incorporate body pose, raw images, face, hand gestures, and gaze as input modalities for the teacher model SocialC3D, enabling it to comprehensively capture uers’ non-verbal social cues. Body pose and raw images are the foundational modalities and prior work has demonstrated their effectiveness and complementarity in understanding body movements (Duan et al., 2022). In contrast, facial expressions, hand gestures, and gaze focus on localized regions of the body, providing fine-grained cues essential for capturing nuanced social signals.
Fig.5 illustrates the performance of SocialC3D and SocialEgoMobile under its supervision with full modality and with each of the three localized modalities ablated. Removing any modality degrades the overall performance of SocialC3D, but the extent of impact varies across social sub-tasks. Specifically, the Attitude and Intent tasks show performance degradation when facial information is removed. Facial orientation offers cues about a user’s attention, while expressions convey emotional states that differentiate social behaviours with similar body movements in early stages, such as ‘throw’ and ‘wave hands’. The Action task is more sensitive to the absence of hand gestures, as detailed hand pose information helps the model to distinguish between fine-grained actions, such as the differences in finger closure for ‘touch’ and ‘handshake’. For the Intent task, the lack of gaze information leads to the greatest drop in accuracy, since gaze direction directly reflects user focus and intention. While facial orientation may offer partial substitutes, it cannot fully replicate the precision of direct gaze signals.
For the student model, 30% of the frames and joints are corrupted in the input. We maintain this corruption ratio to ensure that the findings remain applicable to real-world deployment scenarios. We consider this corruption setting representative, as it strikes a balance between noise and meaningful information. Through knowledge distillation, the impact of missing modalities on the teacher is transferred to the student, as shown in Fig.5. These results confirm that the proposed knowledge distillation framework successfully transfers the integrated multimodal social knowledge from the teacher model to the unimodal student model.
@@
Distillation Methods.  In Table 2, we compare the performance improvements brought by different knowledge distillation methods to SocialEgoMobile. Distillation based on intermediate features outperforms soft label distillation, as the student model in our experiments must learn to integrate information of additional modalities from the teacher’s social representation and compensate for the impact of noisy inputs. Learning only from the teacher’s output distributions is insufficient. Among the feature-based methods, InfoNCE (Oord et al., 2018) achieves the best performance. By maximizing mutual information between the teacher and student model’s social representations, InfoNCE encourages the student to learn semantically aligned features, rather than enforcing exact numerical matching at lower levels. This is suitable for our experiment setting, where the teacher and student operate on different input modalities and the student receives noisy inputs, making low-level alignment impractical. InfoNCE promotes the alignment of high-level semantic information, leading to better noise tolerance and generalization.
@@
Table 3. Comparison of SocialEgoMobile with different architectures when 30% joints and frames are corrupted.
Model Architecture JPL-Social HARPER
SocialEgoMobile 88.10 80.95 75.00 67.31 52.14 57.69
w/ 1 layers GAT 83.33 69.05 67.86 63.46 43.49 48.09
w/ 3 layers GAT 84.52 76.19 69.05 65.38 47.83 53.85
w/ 2 layer GCN (Kipf and Welling, 2017) 85.71 71.43 67.86 59.62 47.83 51.92
w/ 2 layers Bi-LSTM 88.10 82.14 75.00 61.84 52.14 51.92
w/ 1 layer TCN (Lea et al., 2017) 83.33 69.05 65.48 59.62 39.13 48.08
w/ 1 layer Transformer (Vaswani et al., 2017) 77.38 66.67 54.76 55.77 26.08 42.31
@@
Student Model Architecture.  To enhance the robustness and efficiency of processing noisy pose feature sequences, we designed SocialEgoMobile with a two-layer GAT (Veličković et al., 2018) followed by a one-layer Bi-LSTM (Graves and Schmidhuber, 2005). Table 3 presents a comparison of different model configurations under the knowledge distillation with corrupted inputs. In the spatial dimension, a two-layer GAT achieves a good trade-off between expressiveness and stability. Fewer layers limit the receptive field, reducing the ability to capture global graph patterns, while deeper GATs tend to propagate noise more widely. Moreover, GAT’s adaptive attention mechanism enables selective aggregation of useful information from neighbouring joints, offering greater resilience to noisy signals compared to standard Graph Convolutional Networks (GCN) (Kipf and Welling, 2017). In the temporal dimension, a single-layer Bi-LSTM provides sufficient modelling capacity for short feature sequences, while deeper Bi-LSTMs do not result in a steady performance improvement. Compared to Temporal Convolutional Networks (TCN) (Lea et al., 2017) and Transformers (Vaswani et al., 2017), Bi-LSTM is better suited for sequential data with inherent temporal dependencies and benefits from its gating mechanism, which improves robustness to sparsity and temporal discontinuity. Overall, the combination of two GAT layers and one Bi-LSTM in SocialEgoMobile offers an effective balance between expressiveness, robustness, and computational efficiency.